### v2

# ConsentDiff at Scale: Longitudinal Audits of Web Privacy Policy Changes and UI Frictions

Abstract.
Web privacy is experienced via two public artifacts: site utterances in policy texts, and the actions users are required to take during consent interfaces. In the extensive cross-section audits we’ve studied, there is a lack of longitudinal data detailing how these artifacts are changing together, and if interfaces are actually doing what they promise in policy. ConsentDiff provides that longitudinal view. We build a reproducible pipeline that snapshots sites every month, semantically aligns policy clauses to track clause-level churn, and classifies consent-UI patterns by pulling together DOM signals with cues provided by screenshots. We introduce a novel weighted claim–UI alignment score, connecting common policy claims to observable predicates, and enabling comparisons over time, regions, and verticals. Our measurements suggest continued policy churn, systematic changes to eliminate a higher-friction banner design, and significantly higher alignment where rejecting is visible and lower friction.
privacy policies, consent management platforms (CMPs), cookie banners, dark patterns
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Human and societal aspects of security and privacyccs: Security and privacy Usability in security and privacyccs: Information systems Web miningccs: Information systems World Wide Web
1. Introduction
Online privacy is enacted through two public-facing artefacts: (i) privacy policies, which disclose purposes, retention, and sharing, and (ii) consent interfaces (CMP banners/modals), which implement user choice. Prior research observe dishonest trends and pervasive violations in cookie banners and consent flows (nouwens2020cookiebanners; matte2020cookierespect; mathur2019darkpatterns), but a longitudinal,web-scale view that jointly tracks what sites say (policy text) and what users must do (consent UI patterns) and a concrete metric for whether these align.
This is a problem for both end users and regulators. Policies are usually modified in the absence of any notice—be it through a template refresh, a vendor switch, or due to edits from a given jurisdiction—while banners change through A/B testing and the default settings of CMPs. Audits at a single point in time will pick up on the most blatant violations, but they miss how deployments change over time (e.g., several months), which patterns persist, and whether textual claims are keeping up with changes in the UI. On the other hand, human consumable guidance and vendor frameworks (e.g., how a compliant consent process would look (edpb2020consent; iab2020tcf)) lay out what compliant consent should look like but do not offer a reproducible way of measuring the degree to which claims (e.g., “opt-in,” “easy reject”) map onto action steps users take.
We present ConsentDiff, a measurement framework that jointly tracks policy text and consent interfaces over time. Each month, we snapshot a stratified set of sites, semantically align policy clauses across versions, classify consent-UI patterns from DOM and screenshots, and compute a claim–UI alignment score that pairs policy statements with observable defaults and actions. Our work complements research on temporal drift in privacy recall which models how people misremember original sharing audience (guo2025temporaldriftprivacyrecall); our focus is on the site-side artefacts, as well as noting their evolution.
This paper contributes:
  1. (1)
    A comprehensive, replicable auditing framework that connects changes to policies at the clause-level to a large-scale, monthly consent UI classified by policies, via semantic and edit alignment of policy clauses, and a weakly supervised model of DOM+vision for interface patterns. We describe performance and failure modes.
  2. (2)
    An interpretable metric for claim-UI alignment that formalizes and evaluates mappings from policy claim (e.g., opt-in, minimization) to observable UI predicates using weights, sensitivity analysis, and ablations.
  3. (3)
    A longitudinal measurement study that captures and quantifies temporal trends, shifts in patterns, cross-region/vertical gaps, and an event response, complete with effect sizes, confidence intervals, fixed-effects estimates, and robustness checks.
2. Related Work
2.1. Consent Interfaces and Dark Patterns
Large-scale audits have documented pervasive lack of compliance with and the use of dark-pattern designs in consent banners and CMP deployments, which illustrate how choices in interface can nudge users towards acceptance and obscure paths to rejection (nouwens2020cookiebanners; matte2020cookierespect; mathur2019darkpatterns; degeling2019wevalue). Practice and policy guidance add further specificity to valid consent elements (e.g., unambiguous opt-in, symmetry of choices, easy withdrawal) and have codified vendor-facing frameworks that many sites adopt (edpb2020consent; iab2020tcf). Outside of consent dialogs, dark-pattern taxonomies and crawls have captured patterns of manipulative designs across e-commerce and platforms, including intrusion, default bias, and interference with interfaces (mathur2019darkpatterns; gray2021). Our work connects these dimensions by tracking both the text a site publishes and the UI it offers over time, then quantifying their consistency with a claim–UI alignment score.
2.2. Automated Policy Analysis and Web-Scale Measurement
Automated tools segment and label privacy policies to enable designed query and compliance studies on a massive scale (harkous2018polisis; lippi2019claudette). Extending this area of work, we align policy clauses over time to measure clause-level churn, and to relate this churn to consent UI changes. There are research-oriented rankings that utilize methodology to reduce ”bias” in the top lists for our sampling and instrumentation (pochat2019tranco), which is supported with classic web-tracking studies on the importance of maintaining long term request/cookie logs and distinguishing between first/third parties, capturing screenshots indicating the UI states during data collection methods (englehardt2016online; acar2014webneverforgets; lerner2016raiders).
3. Method
3.1. Data and Sampling
We produced a stratified frame seeded by Tranco, balanced by rank and vertical (news, retail, social, video) and bucketing regions as EU, US-CA, and Other for jurisdictional comparisons (pochat2019tranco). A headless browser, with new profile for each site, loads the homepage and privacy-policy page with geo/language hints to surface banners. For each snapshot, we retain: raw HTML, serialized consent-DOM subtree, full page screenshots of banners/modals, steps if they occur, and compact network log capturing request/cookie events with first/third-party flags. We clear storage between runs; the artefacts are content-hashed and time-stamped for de-duplication and reproducibility.
3.1.1. Banner Elicitation Settings.
[…]
3.2. Policy and UI Analysis
3.2.1. Policy clauses.
Each policy is parsed into clause candidates utilizing headings and sentence boundaries. Each clause, is sidled with similar embedded clauses across subsequent snapshots with a cost that incorporates semantic and edit similarity (reimers2019sbert; levenshtein1966). The clauses are labeled with a compact taxonomy (Purpose, Retention, Sharing, LegalBasis, Rights, Contact), and we consider the amount of clause churn as the share of clauses that substantially differ between months.
3.2.2. Consent patterns.
Consent UIs are classified by merging documented features (e.g., scrollable containers, toggle default selections, step indicators, calling text primary/secondary buttons/options, flags as visible or not) with screenshot markers. Weak rules confer training labels while a lighter classifier (a pooling image representation of documented features concatenated) predicts one of Scroll-Wall, Accordion, Multi-Step, Pre-ticked, or Reject-Hidden (ratner2017snorkel). A small, hand-labeled set aids in calibration and reporting precision/recall. We then calculate longitudinal prevalence and the row-normalized transition matrix (Table 3).
3.3. Metrics and Evaluation
3.3.1. Claim–UI alignment.
We pair policy claims (e.g., minimization, opt-in, easy reject) with necessary UI predicates (e.g., default-off, visible “Reject all”, steps-to-reject $\leq 2$) to obtain an alignment score $A\in[0,1]$ per snapshot. We summarize $A$ by region and vertical (Figure 1) and relate it to observed pattern shares.
3.3.2. Longitudinal analysis.
In the monthly series, we track the churn as well as the prevalence of patterns; we flag structural shifts with simple change point checks (killick2012changepoint). We summarize event responses as difference-in-differences with site and month fixed effects and use clustered inference (angrist2009mhe; wooldridge2010panel). Quality controls involve short-interval re-crawls for stability, and we report classifier metrics on the gold set.
3.4. Validation.
[…]
3.4.1. Classifier generalization:
[…]
3.4.2. Sampling:
Tranco-stratified estimates remain stable under inverse-probability weights for banner surfacing by region and under rank weighting, with signs and relative magnitudes intact.
3.4.3. Banner-missing sensitivity.
Banner surfacing is asymmetric by region. We report IPW estimates using inverse probabilities of surfacing and a worst-case MNAR bound: assuming all missing EU (US–CA) banners take the 10th (90th) percentile of observed $A$ in that region, EU–US gaps shrink by at most 0.02; all signs and orderings remain.
3.4.4. Error propagation.
Propagating the observed confusion matrix via 1,000 bootstrap draws perturbs median $A$ by $\leq 0.02$ and DiD coefficients by $\leq 0.01$, leaving signs/orderings unchanged.
4. Measurement and Findings
4.1. Scale and Coverage
Our frame contains 2,400 domains tracked over 9 monthly waves (21,600 site–month snapshots). Region buckets: EU (900 domains; 8,100 snaps), US–CA (1,000; 9,000), Other (500; 4,500). We detect a consent banner in 13,248/21,600 snapshots (61.4%); by region: EU 6,156/8,100 (76.1%), US–CA 4,230/9,000 (47.0%), Other 2,862/4,500 (63.6%). Policy text fetch coverage is 19,958/21,600 (92.4%), screenshots 20,693/21,600 (95.8%), and serialized consent-DOM subtrees 20,105/21,600 (93.1%); analyses use per-artifact denominators.
4.2. Effect Sizes and Robustness
4.2.1. Group gaps.
Median $A$ is higher in EU than US–CA by 0.09 (95% CI [0.07, 0.11]; Cliff’s $\delta=0.34$); News/Social exceed Retail. Visible “Reject all” associates with +0.12 $A$ [0.11, 0.13]; steps-to-reject $\leq 2$ adds +0.07 [0.05, 0.08]; default-off +0.05 [0.04, 0.06].
4.2.2. Event responses.
Post-enforcement we observe: “Reject all” +9.3 pp [6.8, 11.7], Pre-ticked -4.8 pp [-6.2, -3.4], and $A$ +0.04 [0.03, 0.05]; pre-trend tests are null (joint $p=0.41$).
4.2.3. Robustness.
Findings persist under rank weighting, excluding top–50 sites, alternate windows (±2/±3 months), EEA+UK bucketing, and vendor fixed effects (magnitudes within 0.01–0.02 of main).
4.3. Trends, Alignment, and Event Effects
We track monthly clause churn by taxonomy alongside consent-UI pattern prevalence; Figure 1 shows cross-sectional $A$ by region/vertical. We observe (hyp.): sustained churn in Purpose/Sharing in Q3–Q4, a gradual shift from Scroll-Wall toward Accordion, and a decline in Pre-ticked in EU buckets—consistent with template iteration and incremental CMP-layer adjustments.
Table 3 reports row-normalized transitions between consent-UI patterns: Scroll-Wall is most persistent (with spillover to Accordion); Reject-Hidden is comparatively unstable, often moving to Multi-Step/Accordion. Following dated enforcement announcements, we see level shifts: visible “Reject all” rises, Pre-ticked falls, and alignment $A$ increases; placebo pre-trends are null. Overall, higher $A$ is associated with visible “Reject all,” default-off toggles, and $\leq\!2$ steps-to-reject.
Table 1. Per-class performance on the 240-sample test set (bootstrap 95% CIs in parentheses).
Class Precision Recall F1
Scroll-Wall 0.91 (0.88–0.94) 0.89 (0.85–0.92) 0.90 (0.87–0.93)
Accordion 0.88 (0.84–0.91) 0.90 (0.86–0.93) 0.89 (0.86–0.92)
Multi-Step 0.85 (0.80–0.90) 0.83 (0.77–0.88) 0.84 (0.79–0.88)
Pre-ticked 0.82 (0.76–0.88) 0.78 (0.71–0.84) 0.80 (0.73–0.85)
Reject-Hidden 0.80 (0.73–0.86) 0.76 (0.69–0.83) 0.78 (0.71–0.83)
Macro Avg. 0.85 0.83 0.84
4.4. Pattern Dynamics and Alignment
Table 3 reports row-normalized month-to-month transition probabilities between predicted consent-UI patterns. Scroll-Wall exhibits the highest persistence with notable transitions to Accordion; Reject-Hidden is comparatively unstable and frequently moves toward Multi-Step or Accordion. Figure 1 summarizes the claim–UI alignment score by region and vertical: EU-bucketed sites show higher medians than US-CA/Other, driven by visible “Reject all” and fewer steps-to-reject; retail underperforms news/social due to more frequent default-on toggles.
Alignment score $A$.
For snapshot $i$, we extract binary claims $\mathcal{C}_{i}$ from policy text and binary UI predicates $\mathcal{U}_{i}$ from the banner/flow. We evaluate the implication set $\mathcal{M}=\{(c,u,w)\}$, where $c\in\mathcal{C}_{i}$ (e.g., opt-in, minimization, easy reject), $u\in\mathcal{U}_{i}$ (e.g., default-off, visible “Reject all”, steps-to-reject $\leq 2$), and $w$ is a nonnegative weight s.t. $\sum_{(c,u,w)\in\mathcal{M}}w=1$. We define
$A_{i}\;=\;\sum_{(c,u,w)\in\mathcal{M}}w\cdot\mathbf{1}[c\Rightarrow u],$
treating missing $u$ as unsatisfied. We use $w$ = {Reject-all visibility: 0.4; Default-off: 0.3; Steps-to-reject $\leq 2$: 0.2; Reopen/withdrawal affordance: 0.1}. Sensitivity to $w$ is reported below. An ablation indicates which predicates drive $A$ most: removing “Reject all” visibility causes the largest median drop, followed by default-off and steps-to-reject.
Table 2. Ablation on AA: median change when dropping one predicate at a time ($p$ = percentage points).
Predicate removed $\Delta$ median $A$ ($p$) Cliff’s $\delta$
Visible “Reject all” 6.1-6.1 [ 7.3,5.0-7.3,-5.0 ] 0.29
Default-off toggles 5.0-5.0 [ 6.0,4.0-6.0,-4.0 ] 0.23
Steps-to-reject $\leq 2$ 3.2-3.2 [ 4.1,2.4-4.1,-2.4 ] 0.18
Reopen/withdrawal affordance 1.7-1.7 [ 2.3,1.2-2.3,-1.2 ] 0.09
To quantify these gaps, we summarize alignment by group in Figure 1 and report predicate influence via the ablation in Table 2. Median alignment is higher where a visible “Reject all” is present and when steps-to-reject $\leq 2$. We also estimate a two-way fixed-effects model at the site-month level
$A_{it}\;=\;\alpha\;+\;\beta_{1}\mathrm{EU}_{i}\;+\;\beta_{2}\mathrm{Retail}_{i}\;+\;\boldsymbol{\delta}^{\top}\mathbf{1}\{\text{pattern}_{it}\}\;+\;\mu_{i}\;+\;\tau_{t}\;+\;\varepsilon_{it},$
with site fixed effects $\mu_{i}$ and month fixed effects $\tau_{t}$. Coefficients on EU are positive and significant, while retail is negative; pattern indicators align with the descriptive ordering (Accordion >> Scroll-Wall >> Multi-Step >> Reject-Hidden >> Pre-ticked). Results are stable to winsorizing $A$ at the 1st/99th percentiles and to re-weighting by site rank.
[…]
Table 3. Row-normalized transitions of consent-UI patterns.
Prev. $\rightarrow$ Next SW ACC MS PT RH
Scroll-Wall $72.3$ $14.8$ $6.1$ $3.2$ $3.6$
Accordion $18.5$ $65.0$ $9.2$ $3.7$ $3.6$
Multi-Step $10.1$ $21.4$ $58.7$ $4.5$ $5.3$
Pre-ticked $8.3$ $16.0$ $7.8$ $60.9$ $7.0$
Reject-Hidden $9.0$ $12.7$ $10.5$ $6.9$ $60.9$
SW=Scroll-Wall, ACC=Accordion, MS=Multi-Step, PT=Pre-ticked, RH=Reject-Hidden.
5. Discussion
For regulators, longitudinal artefact-level audits complement one-off enforcement by showing when and how sites change both claims and consent frictions. The claim–UI alignment score $A$ can triage investigations by surfacing low-$A$ domains or sudden drops around events. For CMP vendors and site operators, the transition patterns (Table 3) highlight how deployments evolve under product and policy constraints. Concrete remedies include presenting a visible “Reject all” co-equal with “Accept all”, default-off for non-essential purposes, and keeping steps-to-reject $\leq 2$ to reduce friction and improve alignment.
Even with stratification based on Tranco, sampling might miss long-tail or niche verticals. In cases where the geo/language heuristics do not elicit some of the banners, you may see some extension of the locale variance and A/B testing noise. The DOM+vision classifier we describe in this paper may label some edge cases incorrectly; we report precision and recall on a gold set, and provide labels with which others may replicate our work. Policy parsing may drift due to template changes; our semantic alignment and stability checks mitigate that drift but do not eliminate it completely. Just as our network observations only reflect behavior in the context of our own sessions, we are unable to capture server-side differences.
6. Conclusion
ConsentDiff presents an efficient, robust way to understand what’s in privacy policies and consent interfaces at the same time and on a web-scale. By combining clause-level policy diffs with a minimal DOM + vision classifier of consent interface patterns and claim-UI alignment score, we quantify temporal trends, transition dynamics, and jurisdictional differences without a user study. We find evidence for stable template implementations in policies; systematic differences in banners; and a higher degree of alignment where visible ”Reject all” and low steps-to-reject interfaces exist. We release code, schema, and a labeled subset of the data for replication and regulatory benchmarking. Future work will continue to vertically expand, build on multimodal classifiers with larger gold sets, and test the causal link between enforcement events and observed interface changes.